ORTHO-TO-PARA ABUNDANCE RATIO OF CYCLIC-C_{3}H_{2} IN DARK CLOUD CORES

\underline{YUSUKE MORISAWA},TAKAMASA MOMOSE , Division of Chemistry, Graduate School of Science, Kyoto University, Kyoto, 606-8502 JAPAN ; YASUKO KASAI, SMILES Group, Communications Research Laboratory, Nukui-Kita, Koganei, Tokyo, 184-8795,JAPAN; KENTAROU KAWAGUCHI, Department of Chemistry, Fuculty of Science, Okayama University, Okayama, 700-8530 JAPAN.

We have observed four emission lines of cyclic-C_{3}H_{2} in the milimater region with Nobeyama 45m radio telescope toward five cloud cores in TMC-1. From the observed intensities, the density of H_{2} molecules and the fractional abundance of ortho(O) and para(P) of c-C_{3}H_{2} were calculated using the LVG model. The O/P abundance ratio of c-C_{3}H_{2} determined from the calculated abundance is foud to be lower than 3 in the all the observed cloud regions, although both the statistical ratio and the thermal equilibrium ratio at the dust temperature of 10K are predicted to be 3. The deviation of the O/P ratio of c-C_{3}H_{2} from 3 indicates that the O/P ratio of the precursor molecule c-C_{3}H_{3}^{+} is also different from its statistical ratio. Since the precursor molecule c-C_{3}H_{3}^{+} is believed to be produced mainly by the reaction of C_{3}H^{+} + H_{2}, the deviation may be a result of the lower O/P ratio of H_{2} molecules than the statistical value in the observed regions. We will discuss details of the observed data and its analysis.